OpenStudy (anonymous):

If a right triangle has one leg measuring 27 inches and one leg measuring 23 inches, what is the length of the hypotenuse to the nearest inch? 36 14 3 2

OpenStudy (anonymous):

@jessiejamie2

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

you would use pythagorean theorem here a^2+b^2=c^2 so 27^2+23^2=c^2 729+529=c^2 c^2=1258 c=sqrt(1258) c is about 36

OpenStudy (anonymous):

THANKS SO MUCH!! i've been on this test for 2 hours already

OpenStudy (anonymous):

also one more.

OpenStudy (anonymous):

The description of a TV is based upon a diagonal length. If the height of a TV is 24 inches and the width is 32 inches, what is the TV's diagonal length? 3 4 21 40

OpenStudy (anonymous):

again pythagorean theorem since you can make a right traingle from a rectangle 24^2+32^2=c^2 576+1024=c^2 c^2=1600 c=sqrt(1600) c=40

OpenStudy (anonymous):

you actually didnt have to do math

OpenStudy (anonymous):

Really? it's just this confuses me so much it's unbelievable. I"m better at history and language arts :p

OpenStudy (anonymous):

the diagonal length will always be longer than the other sides therefore the only pausible answer from that selection is 40 anyway

OpenStudy (anonymous):

c is always greater than b and a

OpenStudy (anonymous):

lol thanks so much :)

OpenStudy (anonymous):

 I'M GONNA PASS!!